# Large Model Agents: State-of-the-Art, Cooperation Paradigms, Security and Privacy, and Future Trends

### v1

@@ V-A2 Membership Inference Attack (MIA)
MIA refers to inferring whether an individual data sample in the training data of ML models. In the domain of LM agents, MIAs can be further categorized into two types based on the LM training phase: pre-training MIAs and fine-tuning MIAs.
@@
  • Pre-training MIA: The objective of pre-training MIAs is to ascertain whether specific data samples are involved in the training data of pre-trained LMs by analyzing the output generated by LM agents. (i) For LLMs, Mireshghallah et al. [205] propose an innovative MIA that targets Masked Language Models (MLMs) using likelihood ratio hypothesis testing, enhanced by an auxiliary reference MLM. Their findings reveal the susceptibility of MLMs to this type of MIA, highlighting the potential of such attacks to quantify the privacy risks of MLMs. Mattern et al. [206] introduce the neighbourhood MIA, which determines the membership status of target samples by comparing model scores of the given sample with those of synthetic neighbor texts, thus eliminating the need for reference models and enhancing applicability in practical scenarios. Shi et al. [207] present WIKIMIA, a dynamic benchmark for conducting MIAs on pre-training data, using older Wikipedia data as member data and recent Wikipedia data as non-member data. Additionally, they propose a reference-free MIA method named MIN-K% PROB, which computes the average probabilities of outlier tokens to infer membership. (ii) For LVMs, Kong et al. [208] develop an efficient MIA by leveraging proximal initialization. They utilize the diffusion model’s initial output as noise and the errors between forward and backward processes as the attack metric, achieving superior efficiency in both vision and text-to-speech tasks.
@@
  • Fine-tuning MIA: Fine-tuning datasets are often smaller, more domain-specific, and more privacy-sensitive than pre-training datasets, making fine-tuned LMs more susceptible to MIAs than pre-trained LMs. Kandpal et al. [209] introduce a realistic user-level MIA on fine-tuned LMs that utilizes the likelihood ratio test statistic between the fine-tuned LM and a reference model to determine whether a specific user participated in the fine-tuning phase. Mireshghallah et al. [210] conduct an empirical analysis of memorization risks on fine-tuned LMs through MIAs, revealing that fine-tuning the head of the model makes it most susceptible to attacks, while fine-tuning smaller adapters appears to be less vulnerable. Fu et al. [211] propose a self-calibrated probabilistic variation-based MIA, which utilizes the probabilistic variation as a more reliable membership signal, achieving superior performance against overfitting-free fine-tuned LMs.
Figure 23: Illustration of Membership Inference Attack (MIA) to LM agents.
V-A3 Attribute Inference Attack
@@
Attribute inference attacks aim to deduce the presence of specific attributes or characteristics of data samples within the training data of LM agents. For example, such attacks can be exploited to infer the proportion of images with a specific artist style in the training data of a text-to-image agent, potentially leading to privacy breaches for providers of these training images. Pan et al. [212] systematically investigate the privacy risks associated with attribute inference attacks in LMs. Through four diverse case studies, they validate the existence of sensitive attribute inference (e.g., identity, genome, healthcare, and location) within the training data of general-purpose LMs. Wang et al. [213] explore the property existence inference attack against generative models, aiming to determine whether any samples with a target property are contained in the training data. Their study shows that most generative models are susceptible to property existence inference attacks, and they validate this vulnerability in stable diffusion models.
V-A4 Countermeasures to LM Memorization Risks
@@
Existing countermeasures to mitigate memorization risks of LM agents primarily focus on data pre-processing during pre-training and fine-tuning phases. Employing DP techniques and knowledge transfer mechanisms to reduce the LMs’ capacity in memorizing training data during these phases is also a viable approach. Additionally, it is a common strategy to detect and validate privacy disclosure risks of LM agents before deployment.
@@
  • Data sanitization: Data sanitization can effectively mitigate memorization risks by identifying and excluding sensitive information from training data. By replacing sensitive information with meaningless symbols or synthetic data, and removing duplicated sequences, it is possible to defend against privacy attacks that exploit the memorization characteristics of LM agents. Kandpal et al. [214] demonstrate that the rate at which memorized training sequences are regenerated is superlinearly related to the frequency of those sequences in the training data. Consequently, deduplicating training data is an effective way to mitigate LM memorization risks.
@@
  • DP: Existing efforts have validated that adding DP noises to training data and model gradients during pre-training and fine-tuning phases can effectively mitigate the privacy leakages due to LM memorization. Hoory et al. [215] propose a novel differentially private word-piece algorithm, which achieves a trade-off between model performance and privacy preservation capability.
@@
  • Knowledge distillation: Knowledge distillation [123] has been widely adapted as an intuitive technique to preserve privacy. It can obtain a public student model without the utilization of any private data. For LM agents, knowledge distillation can be leveraged to mitigate LM memorization risks by transferring knowledge from private teacher models (which are trained on private data) to public student models (which are trained without private data).
@@
  • Privacy leakage detection & validation: Prior to deploying an LM agent for practical services, it is crucial to mitigate LM memorization risks by detecting and validating the extent of privacy leakage, thereby enabling service providers to modify the model based on validation results. Kim et al. [216] propose ProPILE, an innovative probing tool to evaluate privacy intrusions in LMs. The ProPILE can be employed by LM agent service providers to evaluate the levels of PII leakage for their LMs.
V-B LM intellectual property-related privacy risk
@@
The intellectual property (IP) risks associated with LM agents present two types of privacy risks: LMs-related risks (including LM’s parameters, hyperparameters, and specific training processes), and prompts-related risks (prompts are considered as commodities to generate outputs). The LMs-related information may inherently contain private information, and skilled attackers can further infer private data from this extracted information through carefully crafted privacy attacks. Prompts typically contain user inputs that not only indicate user intent, requirements, and business logic but may also involve confidential information related to the user’s business. We focus on the following two types of IP-related privacy attacks: model stealing attacks and prompt stealing attacks.
@@ V-B1 Model stealing attacks
In model stealing attacks, adversaries aim to extract model information, such as models’ parameters or hyperparameters, by querying models and observing the corresponding responses, subsequently stealing target models without access the original data [217]. Recently, Krishna et al. [218] have demonstrate that language models (e.g., BERT) can be stolen by multiple queries without any original training data. Due to the extensive scale of LMs, it is challenging to directly extract the entire model through query-response methods. Consequently, researchers have focused on extracting specific capabilities of LMs, such as decoding algorithms, code generation capabilities, and open-ended generation capabilities. Naseh et al. [219] demonstrate that an adversary can steal the type and hyperparameters of an LM’s decoding algorithms at a low cost through query APIs. Li et al. [220] investigate the feasibility and effectiveness of model stealing attacks on LMs to extract the specialized code abilities. Jiang et al. [221] propose a novel model stealing attack, which leverages the adversarial distillation to extract knowledge of ChatGPT to a student model through a mere 70k training data, and the student model can achieve comparable open-ended generation capabilities to ChatGPT.
Figure 24: Illustration of LM intellectual property-related privacy risks to LM agents. (a) Model stealing attacks: an adversary maliciously queries the model with multiple similar questions to obtain a series of response pairs. This allows them to steal the LM, leading to privacy breaches or the creation of competing products. (b) Prompt stealing attacks: attributes of the original prompt are determined using a subject generator and a modifier detector, and the reverse prompt is reconstructed, resulting in privacy exposure.
@@ V-B2 Prompt stealing attacks
With the advancement of LM agent services, high-quality prompts designed to generate expected content have acquired substantial commercial value. These prompts can be traded on various prompt marketplaces, such as PromptSea [^11] https://www.promptsea.io/ and PromptBase [^12] https://promptbase.com/. Consequently, a new privacy attack called prompt stealing attack has emerged, where an adversary aims to infer the original prompt from the generated content. This attack is analogous to the model inversion attack in traditional ML, which involves reconstructing the input based on the output of an ML model [222]. Shen et al.[223] conduct the first study on prompt stealing attack in text-to-image generation models, and propose an effective prompt stealing attack method named PromptStealer. The PromptStealer utilizes a subject generator to infer the subject and a modifier detector to identify the modifiers within the generated image. Sha et al.[224] extend the prompt stealing attack to LLMs, using a parameter extractor to determine the properties of original prompts and a prompt reconstructor to generate reversed prompts.
V-B3 Countermeasures to Model & Prompt Stealing Attacks
@@
  • Model watermarking: Model watermarking is an innovative technique in protecting IP rights and ensuring accountability for LM agents. By embedding watermarks to target LMs, the ownership of LMs can be authenticated by verifying the watermarks, thereby preventing unauthorized use or infringement. Kirchenbauer et al. [225] propose a watermarking algorithm utilizing a randomized set of “green” tokens during the text generation process, where the model watermark is verified by a statistical test with interpretable p-values.
@@
  • Blockchain: Blockchain can be employed as a transparent platform to verify IP rights due to its inherent immutability and traceability [101]. The owner of LMs can record the develop logs, version information, and hash values of LMs’ parameters on blockchain, ensuring the authenticity and completeness of the recorded information. Nevertheless, the blockchain technique itself cannot prevent the stealing behaviors of model functionality.
@@
  • Adversarial perturbations: Transforming the generated content into adversarial examples by adding optimized perturbations is an effective method to prevent prompt-stealing attacks while maintaining the quality of the generated content. Shen et al. [223] propose an intuitive defense mechanism named PromptShield, which employs the adversarial example technique to add a negligible perturbation on generated images, thereby defending against their proposed prompt stealing attack PromptStealer. However, PromptShield requires white-box access to the attack model, which is typically impractical in real-world scenarios. Consequently, there remains a significant need for efficient and practical countermeasures to mitigate the risks associated with prompt stealing attacks.
V-C Other Privacy Threats to LM Agents
@@
  • Sensitive query attack: In LM agent services, the LM may memorize sensitive personal or organizational information in user queries, resulting in potential privacy leakages. For example, Samsung employees leveraged ChatGPT for code auditing without processing the confidential information in Apr. 2023, inadvertently exposing the company’s commercial secrets including source code of the new program [226].
@@
  • Privacy leakage in multi-agent interactions: LM agent services typically necessitate seamless collaboration of multiple LM agents to address complex user queries, where each agent is tasked with solving particular sub-problems of the queries. Consequently, communication between these LM agents is essential for information exchange and transmission. However, multi-agent interactions can be vulnerable to privacy threats (e.g., eavesdropping, compromised agent, and man-in-the-middle attacks), leading to potential user privacy breaches. Since interactions in LM agent services typically occur through natural language, traditional methods such as homomorphic encryption and secure multi-party computation struggle to effectively safeguard the privacy of these interactions. It remains a challenge to design new strategies tailored to these specific vulnerabilities to preserve privacy in multi-agent interactions.
@@ V-D Summary and Lessons Learned
There are primarily two types of privacy threats to LM agents: LM memorization risk, and LM IP-related privacy risk. Generally, data extraction attacks, MIAs, and attribute inference attacks are three main privacy threats stemmed from LM memorization risks. Besides, model stealing attacks and prompt stealing attacks are two typical LM IP-related privacy risks. Other privacy threats to LM agents include sensitive query attacks and privacy leakage in multi-agent interactions. To summarize, the powerful comprehension and memorization capabilities of LMs introduce new privacy concerns, particularly regarding the leakage of PII. Meanwhile, the interaction modes of LM agents have endowed prompts with commercial value, highlighting the importance of intellectual property rights associated with them. Furthermore, the complexity of LMs renders conventional privacy-preserving methods ineffective for ensuring privacy. Therefore, to comprehensively safeguard privacy within LM agent systems, researchers should develop effective and innovative privacy protection techniques tailed for LM agents. Additionally, it is imperative for governments and authoritative organizations to advance legislation process related to privacy breaches and intellectual property of LM agent services.

### v2

@@ V-A2 Membership Inference Attacks (MIAs)
MIA refers to inferring whether an individual data sample in the training data of AI models. In the domain of LM agents, MIAs can be further categorized into two types based on the LM training phase: pre-training MIA and fine-tuning MIA.
@@
  • Pre-training MIA: As shown in Fig. 30, the objective of pre-training MIAs is to ascertain whether specific data samples are involved in the training data of pretrained LMs by analyzing the output generated by LM agents. (i) For LLMs, Mireshghallah et al. [187] propose an innovative MIA that targets masked language models (MLMs) using likelihood ratio hypothesis testing, enhanced by an auxiliary reference MLM. Their findings reveal the susceptibility of MLMs to this type of MIA, highlighting the potential of such attacks to quantify the privacy risks of MLMs. (ii) For LVMs, Kong et al. [188] develop an efficient MIA by leveraging proximal initialization. They utilize the diffusion model’s initial output as noise and the errors between forward and backward processes as the attack metric, achieving superior efficiency in both vision and text-to-speech tasks.
@@
  • Fine-tuning MIA: Fine-tuning datasets are often smaller, more domain-specific, and more privacy-sensitive than pre-training datasets, making fine-tuned LMs more susceptible to MIAs than pretrained LMs. Kandpal et al. [189] introduce a realistic user-level MIA on fine-tuned LMs that utilizes the statistic of likelihood ratio test between the fine-tuned LM and a reference model to determine whether a specific user participated in the fine-tuning phase. Mireshghallah et al. [190] conduct an empirical analysis of memorization risks on fine-tuned LMs through MIAs, revealing that fine-tuning the head of the model makes it most susceptible to attacks, while fine-tuning smaller adapters appears to be less vulnerable. Fu et al. [191] propose a self-calibrated probabilistic variation-based MIA, which utilizes the probabilistic variation as a more reliable membership signal, achieving superior performance against overfitting-free fine-tuned LMs.
V-A3 Attribute Inference Attacks
@@
Attribute inference attacks aim to deduce the presence of specific attributes or characteristics of data samples within the training data of LM agents. For instance, such attacks can be exploited to infer the proportion of images with a specific artist style in the training data of a text-to-image agent, potentially leading to privacy breaches for providers of these training images. Pan et al. [192] systematically investigate the privacy risks associated with attribute inference attacks in LLMs. Through four diverse case studies, they validate the existence of sensitive attribute inference (e.g., identity, genome, healthcare, and location) within the training data of general-purpose LLMs. Wang et al. [193] explore the property existence inference attack against generative models, aiming to determine whether any samples with a target property are contained in the training data. Their study shows that most generative models are susceptible to property existence inference attacks, and they validate this vulnerability in stable diffusion models.
V-A4 Countermeasures to LM Memorization Risks in LM Agent Networks
@@
Existing countermeasures to mitigate memorization risks of LM agents primarily focus on data pre-processing during pre-training and fine-tuning phases. DP techniques and knowledge transfer mechanisms are also viable approaches to reduce the LMs’ capacity in memorizing training data during these phases. Additionally, it is a common strategy to detect and validate privacy disclosure risks of LM agents before deployment.
@@
  • Data sanitization can effectively mitigate memorization risks by identifying and excluding sensitive information from training data. By replacing sensitive information with meaningless symbols or synthetic data, and removing duplicated sequences, it is possible to defend against privacy threats that exploit the memorization characteristics of LM agents. Kandpal et al. [194] demonstrate that the rate at which memorized training sequences are regenerated is superlinearly related to the frequency of those sequences in the training data. Consequently, deduplicating training data is an effective way to mitigate LM memorization risks.
@@
  • Adding DP noises to training data and model gradients during pre-training and fine-tuning phases can effectively mitigate privacy leakages due to LM memorization. For instance, Hoory et al. [195] propose a novel differentially private word-piece algorithm, which achieves a trade-off between model performance and privacy preservation capability.
@@
  • Knowledge distillation [113] has been widely adopted as an intuitive technique to preserve privacy. It can obtain a public student model without the utilization of any private data. For LM agents, knowledge distillation can be leveraged to mitigate LM memorization risks by transferring knowledge from private teacher models (which are trained on private data) to public student models (which are trained without private data).
@@
  • Privacy leakage detection and assessment can effectively mitigate LM memorization risks before deploying an LM agent for practical services, thereby enabling service providers to modify the model based on assessment results. For instance, Kim et al. [196] propose ProPILE, an innovative probing tool to evaluate privacy intrusions in LLMs. ProPILE can be employed by service providers to evaluate the levels of PII leakage of their LLM agents.
V-B LM Stealing & Prompt Stealing Attacks to LM Agent Networks
@@
In LM agent networks, LM stealing risks (including LM’s parameters, hyperparameters, and specific training processes) and prompt stealing risks (prompts are considered as commodities to generate outputs) are two types of intellectual property (IP)-related privacy threats to LM agents, as illustrated in Fig. 31. LM-related information may inherently contain private information, and smart attackers can further infer private data from the extracted information through carefully crafted privacy threats. Besides, prompts typically contain user inputs that not only indicate user intent, and requirements, but may also involve confidential business logic.
Figure 31: Illustration of IP-related privacy risks to LM agents: (a) model stealing attack; (b) prompt stealing attack.
@@ V-B1 Model Stealing Attacks
In model stealing attacks, adversaries aim to extract model information, such as model parameters or hyperparameters, by querying models and observing the corresponding responses, subsequently stealing target models without access the original data [197]. Recently, Krishna et al. [198] demonstrate that language models (e.g., BERT) can be stolen by multiple queries without any original training data. Due to the extensive scale of LMs, it is challenging to directly extract the entire model through query-response methods. Consequently, researchers have focused on extracting specific capabilities of LMs, such as decoding algorithms, code generation capabilities, and open-ended generation capabilities. Naseh et al. [199] demonstrate that adversaries can steal the type and hyperparameters of an LM’s decoding algorithms at a low cost through query APIs. Li et al. [200] investigate the feasibility and effectiveness of model stealing attacks on LLMs to extract the specialized code abilities. Jiang et al. [201] propose a novel model stealing attack, which leverages the adversarial distillation to extract knowledge of ChatGPT to a student model through a mere 70k training data, and the student model can achieve comparable open-ended generation capabilities to ChatGPT.
@@ V-B2 Prompt Stealing Attacks
With the advancement of LM agent services, high-quality prompts designed to generate expected content have acquired substantial commercial value. These prompts can be traded on various prompt marketplaces, such as PromptSea [^12] https://www.promptsea.io/ and PromptBase [^13] https://promptbase.com/. Consequently, a new privacy threat called prompt stealing attack has emerged, where an adversary aims to infer the original prompt from the generated content. This attack is analogous to the model inversion attack in traditional AI, which involves reconstructing the input based on the output of an AI model [202]. Shen et al.[203] conduct the first study on prompt stealing attack in text-to-image generation models, and propose an effective prompt stealing attack method named PromptStealer. PromptStealer utilizes a subject generator to infer the subject and a modifier detector to identify the modifiers within the generated image. Sha et al.[204] extend the prompt stealing attack to LLMs, using a parameter extractor to determine the properties of original prompts and a prompt reconstructor to generate reversed prompts. Hui et al.[205] propose an innovative closed-box prompt extraction framework, called PLEAK, which employs incremental search techniques and post-processing methods to optimize adversarial queries for the purpose of extracting system prompts from LLM applications. The effectiveness of PLEAK is demonstrated through its evaluation on real-world applications hosted on the Poe platform  [^14] https://poe.com.
TABLE VIII: Summary of Key Literature on Privacy Threat Countermeasures to LM Agent Networks
Ref.
Privacy
Threat
$\star$ Purpose
$\bullet$ Advantages
$\circ$ Limitations
$\dagger$ Evaluation Metrics
Defense
Methods
[194]
LM memorization
threats
$\star$ Reduce the privacy leakage risk due to memorization of training data in language models
$\bullet$ Privacy enhancement, performance preservation, and simplicity
$\circ$ Lack of suitable duplication match mechanisms, lack of evaluation on diverse data domains
$\dagger$ AUROC(Area under ROC curve), TPR
Data
sanitization
[195]
LM memorization
threats
$\star$ Protect sensitive information in pre-trained language models’ training data from leakage
$\bullet$ Privacy and performance preservation, medical domain adaptability
$\circ$ Lack of evaluation on large-scale models, lack of evaluation on different pre-training data
$\dagger$ F1 scores
DP
[113]
LM memorization
threats
$\star$ Boost performance of small language models, reducing computation costs and privacy risks
$\bullet$ Lower computational overheads, privacy preservation, and high sample efficiency
$\circ$ Lack of exploration on LLMs, heavy dependency on external knowledge base
$\dagger$ Accuracy
Knowledge
distillation
[196]
LM memorization
threats
$\star$ Assess PII leakage risk of LLMs for data owners
$\bullet$ User autonomy, high scalability and flexibility
$\circ$ Heavy dependency on assess dataset, potential misuse risks
$\dagger$ Percentage of exact match results, reconstruction likelihood
Privacy leakage
assessment
[206]
LM stealing
attacks
$\star$ Detect and trace model-generated text by embedding watermarks to generated content by LLMs
$\bullet$ Model-agnostic detection, robust watermarks, and cost-saving
$\circ$ Cannot defend semantic-based attacks, lack of extensive theoretical analysis
$\dagger$ z-score, FPR, FNR, AUROC
Model
watermarking
[203]
Prompt stealing
attacks
$\star$ Mitigate proposed prompt stealing attacks utilizing adversarial examples
$\bullet$ High effectiveness and utility, high feasibility in practical scenarios
$\circ$ Strong assumption, weak transfer defense ability, and vulnerability to adaptive attacks
$\dagger$ Cosine similarity, Jaccard similarity, mean squared error (MSE)
Adversarial
examples
V-B3 Countermeasures to Model & Prompt Stealing Attacks in LM Agent Networks
@@
  • Model watermarking is an innovative technique to protect IP rights and ensure accountability for LM agents. By embedding watermarks to target LMs, the ownership of LMs can be authenticated by verifying the watermarks, thereby preventing unauthorized use or infringement. For instance, Kirchenbauer et al. [206] propose a watermarking algorithm utilizing a randomized set of “green” tokens during the text generation process, where the model watermark is verified by a statistical test with interpretable p-values.
@@
  • Blockchain can be employed as a transparent platform to verify IP rights due to its inherent immutability and traceability [207]. The owner of LMs can record the develop logs, version information, and hash values of LM parameters on blockchain, ensuring the authenticity and completeness of the recorded information. Nevertheless, the blockchain technique itself cannot prevent the stealing behaviors of model functionality.
@@
  • Adversarial examples is an effective method to prevent prompt-stealing attacks while maintaining the quality of the generated content by adding optimized perturbations to generated content. Shen et al. [203] propose an intuitive defense mechanism named PromptShield, which employs the adversarial example technique to add a negligible perturbation on generated images, thereby defending against the proposed prompt stealing attack PromptStealer. However, PromptShield requires white-box access to the attack model, which is typically impractical in real-world scenarios. Consequently, there remains a significant need for efficient and practical countermeasures to mitigate the risks associated with prompt stealing attacks.
V-C Other Privacy Threats to LM Agent Networks
@@
  • Sensitive query attacks: In LM agent services, the LM may memorize sensitive personal or organizational information in user queries, resulting in potential privacy leakages. For instance, Samsung employees leveraged ChatGPT for code auditing without processing the confidential information in Apr. 2023, inadvertently exposing the company’s commercial secrets including source code of the new program.
@@
  • Privacy leakage in multi-agent interactions: LM agent services typically necessitate seamless collaboration among multiple LM agents to address complex user queries, where each agent is tasked with solving particular sub-problems of the queries. Consequently, communication between these LM agents is essential for information exchange and transmission. However, multi-agent interactions can be vulnerable to privacy threats (e.g., eavesdropping, compromised agent, and man-in-the-middle attacks), leading to potential user privacy breaches. Since interactions in LM agent services typically occur through natural language and semantic meanings, traditional methods such as homomorphic encryption and secure multi-party computation struggle to effectively safeguard the privacy of these semantically rich interactions. It remains a challenge to design new strategies tailored to these specific vulnerabilities to preserve privacy during multi-agent interactions.
@@ V-D Summary and Lessons Learned
There are primarily two types of privacy threats to LM agents: LM memorization threats, and IP-related privacy threats. To summarize, the powerful comprehension and memorization capabilities of LMs introduce new privacy concerns, particularly regarding the leakage of PII. Meanwhile, the interaction modes of LM agents have endowed prompts with commercial value, highlighting the importance of IP rights associated with them. Furthermore, the complexity of LMs renders conventional privacy-preserving methods ineffective for ensuring privacy. Therefore, to comprehensively safeguard privacy within LM agent systems, researchers should develop effective and innovative privacy protection techniques tailed for connected LM agents. Additionally, it is imperative for governments and authoritative organizations to advance legislation process related to privacy breaches and IP of LM agent services.